[ The length of the rectangle is 8 centimeters less than twice the width. The area of the rectangle is 24 centimeters squared. Determine the dimensions of the rectangle in centimeters ]
Answer by TimothyLamb(4379)

L = 2w - 8
Lw = 24
---
Lw = 24
(2w - 8)w = 24
2w^2 - 8w - 24 = 0
---
---
the above quadratic equation is in standard form, with a=2, b=-8, and c=-24
---
to solve the quadratic equation, plug this:
2 -8 -24
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the roots of the quadratic are:
w = 6
w = -2
---
the negative root doesn't make sense for a length, so use the positive root
---
Answer:
w = 6 cm
L = 4 cm
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php